# Wall crossing and the Fourier–Mukai transform for Grassmann flops

### v1

@@
Abstract.
@@
1. Introduction
The connection between birational geometry and quantum cohomology has been a source of great activity and interest in Gromov–Witten theory since the early development of the theory. As first described in Ruan’s crepant resolution conjecture in [Rua99] and later extended to more general crepant transformations in [LR01, CIT09, CR13, LLW10], it is expected that for a general crepant rational map $f:X_{-}\dashrightarrow X_{+}$, the quantum cohomology of $X_{-}$ should be equivalent to that of $X_{+}$ in a specific sense.
@@ 1. Introduction
such that $f_{-}^{*}(K_{X_{-}})=f_{+}^{*}(K_{X_{+}})$. The crepant transformation conjecture predicts that generating functions of genus zero Gromov–Witten invariants of $X_{-}$ and $X_{+}$ are equal after analytic continuation. Versions of the conjecture have now been proven in many specific instances, and for certain classes of crepant transformations (see e.g. [LR01, BG09, CIJ18, ILLW12]).
One can further refine the crepant transformation conjecture to take into account Iritani’s integral structure (see [Iri09]) or the rational structure of Katzarkov–Kontsevich–Pantev (see [KKP08]). With respect to this structure, the correspondence should be compatible with the Fourier–Mukai transform
$\mathbb{FM}={f_{+}}_{*}f_{-}^{*}:K^{0}(X_{-})\to K^{0}(X_{+}).$
@@
1.1. Grassmann flops
A rich source of examples of crepant transformations as in (1.0.1) arises from variation of GIT. Namely, suppose a reductive group $G$ acts on a variety $V$. Then we can define two GIT quotients
$X_{\pm}=[V\sslash_{\pm}G],$
where $+/-:G\to\mathbb{C}^{*}$ are characters lying in adjacent maximal chambers with respect to the wall and chamber structure on $\operatorname{Hom}(G,\mathbb{C}^{*})\otimes_{\mathbb{Z}}\mathbb{R}$. Under certain numerical conditions, the rational map $X_{+}\dashrightarrow X_{-}$ will be crepant. In this context, the relation between the Gromov–Witten theory of $X_{+}$ and $X_{-}$ is known as wall crossing, as the variation of GIT between the two varieties involves crossing a codimension-one wall in the $G$-ample cone.
@@ 1.1. Grassmann flops
The setup is as follows: given a smooth projective variety $B$, let $F=\oplus_{i=1}^{n}M_{i}$ and $E=\oplus_{j=1}^{n}L_{j}$ be sums of line bundles on $B$. For $r<n$, define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}om(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}om(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
There is a natural (fiberwise) action of $G=\operatorname{GL}_{r}$ on $V$. Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$ respectively. Then
$X_{+}=$ $\operatorname{tot}(S_{+}\otimes p^{*}F^{\vee})\to\operatorname{Gr}(r,E)$
$X_{-}=$ $\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}(r,F^{\vee})$
where $p:\operatorname{Gr}(r,E)\to B$ (resp. $p:\operatorname{Gr}(r,F^{\vee})\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. We will work equivariantly with respect to the action of $\mathbb{T}=(\mathbb{C}^{*})^{2n}$ induced by the natural action on $E\oplus F$.
@@
We give an example of a relative Grassmann flop realizing a variation of GIT of quiver varieties.
Fix integers $0<r<k<n$ and let $B$ be the Grassmannian $\operatorname{Gr}(k,n).$ Let $F\to B$ be the trivial rank $k$ bundle $\mathcal{O}_{B}\otimes\mathbb{C}^{k}$ and let $E\to B$ be the rank $k$ tautological bundle on $B$. In this case, $X_{+}$ is
$X_{+}=\operatorname{tot}(\hat{S}_{r}^{\oplus k})$
@@
$X_{-}=\operatorname{tot}(\bar{S}_{r}^{\vee}\otimes S_{k}).$
@@
$\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})\times\operatorname{Hom}(\mathbb{C}^{r},\mathbb{C}^{k})\times\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$
by the natural right action of $\operatorname{GL}_{r}\times\operatorname{GL}_{k}$ where $\operatorname{GL}_{r}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{n})$ and $\operatorname{GL}_{k}$ acts trivially on $\operatorname{Hom}(\mathbb{C}^{k},\mathbb{C}^{r})$. Then $X_{+}$ and $X_{-}$ from above are GIT quotients of this representation associated to the characters
$+:(g_{1},g_{2})\mapsto\det(g_{1})\cdot\det(g_{2})$
and
$-:(g_{1},g_{2})\mapsto\det(g_{1})^{-1}\cdot\det(g_{2})$
@@
1.2. Statement of theorem
With $X_{\pm}$ as above, consider the infinite dimensional (symplectic) vector space
$\mathcal{H}_{X_{\pm}}=H^{*}_{\mathbb{T}}(X_{\pm})(\!(z^{-1})\!)[\![Q,q_{\pm}]\!].$
@@ 1.2. Statement of theorem
Using the relative quasimap theory of [Oh21], one can define $I$-functions
$I_{X_{\pm}}(Q,q_{\pm},z)$
lying in $\mathcal{H}_{X_{\pm}}$. These functions are shown to lie on Givental’s Lagrangian cone $\mathcal{L}_{X_{\pm}}\subset\mathcal{H}_{X_{\pm}}$ for the respective spaces and may therefore be expressed as generating functions of genus-zero Gromov–Witten invariants of $X_{\pm}$. See Section 3 for details.
One can then extend $\mathcal{H}_{X_{\pm}}$ by ~$\widetilde{\mathcal{H}}_{X_{\pm}}=\mathcal{H}_{X_{\pm}}[\log z][z^{-1/2}]$, and following Iritani in [Iri09], define the map
$\Psi:K^{0}_{\mathbb{T}}(X)$$\to\widetilde{\mathcal{H}}_{X}$
$A$ $\mapsto z^{-\mu}z^{\rho}\left(\hat{\Gamma}_{X}\cup(2\pi\mathbbm{i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\right)$
where $\mathbbm{i}=\sqrt{-1}$, $\mu$ is the grading operator, $\rho=c_{1}^{\mathbb{T}}(X_{\pm})$, and $\hat{\Gamma}_{X}$ is the Gamma class—a characteristic class associated with the Gamma function $\Gamma(1+z)$.
We prove the following:
@@
There exists a linear symplectic isomorphism
$\mathbb{U}:\widetilde{\mathcal{H}}_{X_{-}}\to\widetilde{\mathcal{H}}_{X_{+}}$
such that the following diagram commutes:
(1.2.1) ${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\mathbb{U}}$
@@
Theorem 1.2 generalizes previous known cases of the crepant transformation conjecture in two directions. First, it is to our knowledge one of the first examples of a wall crossing result for a variation of GIT quotients by a non-abelian group—in this case $\operatorname{GL}_{r}$. Second, it provides an example of the conjecture for fiber bundles over an arbitrary smooth base $B$. Thus our result is new even in the special case of the quotient by $\operatorname{GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over $B$.
1.3. Strategy of proof
@@
$I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,\pm},\ldots,q_{r,\pm},z)$
@@
at $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}.$ Here $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ is a twist of the $I$-function for the associated abelian quotient by the root bundle for $G$, which we denote by $\mathfrak{g}/\mathfrak{t}$. See Section 3 for details.
@@
The technical heart of the proof relies on computing the analytic continuation of $I_{X_{+}}$ along the path $\hat{\gamma}$ and using this to determine $\mathbb{U}$. It is not clear how to compute $\widetilde{I_{X_{+}}}$ directly. Instead we compute the analytic continuation of $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along each of the variables $q_{k,+}$ independently. That is to say, we analytically continue $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\ldots,q_{r,+},z)$ along a concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$, where each $\gamma_{k}$ leaves the variables $q_{i,+}$ constant for $i\neq k$, and goes from 00 to $\infty$ in the $q_{k,+}$ direction. We then homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ to a path $\hat{\gamma}$ which lies entirely in the locus $q_{1,+}=\cdots=q_{r,+},$ thereby obtaining $\widetilde{I_{X_{+}}}$.
@@
Finally, we must verify that the linear map $\mathbb{U}$ identifying $I_{X_{-}}$ with $\widetilde{I_{X_{+}}}$ is compatible with the Fourier–Mukai transform. This involves an explicit computation of $\mathbb{FM}$ on generators of $K^{0}_{\mathbb{T}}(X_{-})$. The proof that the map $\mathbb{U}$ in (1.2.1) satisfies (1.2.2) then involves a nontrivial identity of anti-symmetric functions (Lemma 5.2).
1.4. Relation to other work
@@
The structure of the proof follows that given in [CIT09, CR10, CIR14, CIJ18]. In particular the formulation of the correspondence and the proof via localization is closely related to [CIJ18]. The new ingredients in the current article are the homotopy of the path of analytic continuation, the generalization to the relative setting, and the combinatorial identity verifying agreement with $\mathbb{FM}$, which does not appear in the case of abelian variation of GIT.
@@
In the final stages of writing this paper, we learned of a related result by Lutz–Shafi–Webb (see [LSW]). While both papers use the idea of homotoping the path of analytic continuation, there are significant differences in the two results and their proofs. The most important such difference is in the analysis of the symplectic transformation $\mathbb{U}$. In this paper, we compare the map $\mathbb{U}$ to the Fourier–Mukai transform $\mathbb{FM}$. In [LSW], they instead compare $\mathbb{U}$ with the symplectic transformation of the associated abelian wall crossing. This leads to two different proofs that the map $\mathbb{U}$ is a symplectic isomorphism and has a well-defined non-equivariant limit. See Remark 5.5 for an explicit comparison of approaches.
The two papers are in fact complementary, in that our respective results may be combined to show that the abelian/non-abelian correspondence is compatible with the Fourier–Mukai transform in $K$-theory. Such a compatibility is not a priori obvious, as it is not clear how best to relate the kernel $\tilde{X}$ of the Fourier–Mukai transform of the Grassmann flop to the kernel of Fourier–Mukai transform of the associated abelian wall crossing. We hope to revisit this issue in the future to obtain a more direct and conceptual understanding of this compatibility.
1.5. Acknowledgements
We would like to thank Tom Coates, Wendelin Lutz, Jeongseok Oh, Yongbin Ruan, Ed Segal, Qaasim Shafi, and Rachel Webb for the useful discussions.
N. Priddis is supported by the Simons Foundation Travel Grant 586691. M. Shoemaker is supported by the Simons Foundation Travel Grant 958189. Y. Wen is supported by a KIAS Individual Grant (MG083902) at the Korea Institute for Advanced Study.
2. Geometric setup

### v2

@@
[…]
Wall Crossing and the Fourier–Mukai Transform for Grassmann Flops
[…]
Wall Crossing and the Fourier–Mukai Transform
for Grassmann Flops
[…]
Nathan PRIDDIS a, Mark SHOEMAKER b and Yaoxiong WEN c
[…]
N. Priddis, M. Shoemaker and Y. Wen
[…]
b) Department of Mathematics, Colorado State University,
[…]
Received April 26, 2024, in final form January 27, 2025; Published online February 06, 2025
[…]
@@
[…]
Fourier–Mukai; Grassmannian flops; wall-crossing; Gromov–Witten theory; variation of GIT
[…]
14N35; 14E16; 53D45
1 Introduction
The connection between birational geometry and quantum cohomology has been a source of great activity and interest in Gromov–Witten theory since the early development of the theory. As first described in Ruan’s crepant resolution conjecture in [31] and later extended to more general crepant transformations in [16, 18, 25, 27], it is expected that for a general crepant rational map $f\colon X_{-}\dashrightarrow X_{+}$, the quantum cohomology of $X_{-}$ should be equivalent to that of $X_{+}$ in a specific sense.
@@ 1 Introduction
such that $f_{-}^{*}(K_{X_{-}})=f_{+}^{*}\bigl{(}K_{X_{+}}\bigr{)}$. The crepant transformation conjecture predicts that generating functions of genus zero Gromov–Witten invariants of $X_{-}$ and $X_{+}$ are equal after analytic continuation. Versions of the conjecture have now been proven in many specific instances, and for certain classes of crepant transformations (see, e.g., [8, 14, 23, 27]).
One can further refine the crepant transformation conjecture to take into account Iritani’s integral structure (see [22]) or the rational structure of Katzarkov–Kontsevich–Pantev (see [24]). With respect to this structure, the correspondence should be compatible with the Fourier–Mukai transform
$\mathbb{FM}={f_{+}}_{*}f_{-}^{*}\colon\ K^{0}(X_{-})\to K^{0}\bigl{(}X_{+}\bigr{)}.$
@@
1.1 Grassmann flops
A rich source of examples of crepant transformations as in (1.1) arises from variation of GIT. Namely, suppose a reductive group $G$ acts on a variety $V$. Then we can define two GIT quotients $X_{\pm}=[V\sslash_{\pm}G]$, where $+/-\colon G\to\mathbb{C}^{*}$ are characters lying in adjacent maximal chambers with respect to the wall and chamber structure on $\operatorname{Hom}(G,\mathbb{C}^{*})\otimes_{\mathbb{Z}}\mathbb{R}$ [^1] For those unfamiliar with the notation, by $[V\sslash_{\pm}G]$ we mean the quotient stack $[V^{\rm ss}(\pm)/G]$. Under certain numerical conditions, the rational map $X_{+}\dashrightarrow X_{-}$ will be crepant. In this context, the relation between the Gromov–Witten theory of $X_{+}$ and $X_{-}$ is known as wall crossing, as the variation of GIT between the two varieties involves crossing a codimension-one wall in the $G$-ample cone.
@@ 1.1 Grassmann flops
The setup is as follows: given a smooth projective variety $B$, let $F=\bigoplus_{i=1}^{n}M_{i}$ and $E=\bigoplus_{j=1}^{n}L_{j}$ be sums of line bundles on $B$. For $r<n$, define the vector bundle $V\to B$ to be the total space of
$\mathcal{H}{\rm om}(F,\mathcal{O}_{B}\otimes\mathbb{C}^{r})\times\mathcal{H}{\rm om}(\mathcal{O}_{B}\otimes\mathbb{C}^{r},E).$
There is a natural (fiberwise) action of $G={\rm GL}_{r}$ on $V$. Let $X_{+}$ and $X_{-}$ denote the GIT quotients $V\sslash_{\pm}G$ associated to the characters $\det^{+1}$ and $\det^{-1}$, respectively. Then
$X_{+}=\operatorname{tot}\bigl{(}S_{+}\otimes p^{*}F^{\vee}\bigr{)}\to\operatorname{Gr}(r,E),\qquad X_{-}=\operatorname{tot}(S_{-}\otimes p^{*}E)\to\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)},$
where $p\colon\operatorname{Gr}(r,E)\to B$ (resp. $p\colon\operatorname{Gr}\bigl{(}r,F^{\vee}\bigr{)}\to B$) denotes the relative Grassmannian of $r$-planes in $E$ (resp. $F^{\vee}$) and $S_{\pm}$ denotes the rank-$r$ tautological bundle on the associated Grassmann bundle. We will work equivariantly with respect to the action of $\mathbb{T}=(\mathbb{C}^{*})^{2n}$ induced by the natural action on $E\oplus F$.
@@
We give an example of a relative Grassmann flop realizing a variation of GIT of quiver varieties. Fix integers $0<r<k<n$ and let $B$ be the Grassmannian $\operatorname{Gr}(k,n).$ Let ${F\to B}$ be the trivial rank $k$ bundle $\mathcal{O}_{B}\otimes\mathbb{C}^{k}$ and let $E\to B$ be the rank $k$ tautological bundle on $B$. In this case,
$X_{+}=\operatorname{tot}\bigl{(}\hat{S}_{r}^{\oplus k}\bigr{)},$
@@
$X_{-}=\operatorname{tot}\bigl{(}\bar{S}_{r}^{\vee}\otimes S_{k}\bigr{)}.$
@@
$\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{r}\bigr{)}\times\operatorname{Hom}\bigl{(}\mathbb{C}^{r},\mathbb{C}^{k}\bigr{)}\times\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{n}\bigr{)}$
by the natural right action of ${\rm GL}_{r}\times{\rm GL}_{k}$ where ${\rm GL}_{r}$ acts trivially on $\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{n}\bigr{)}$ and ${\rm GL}_{k}$ acts trivially on $\operatorname{Hom}\bigl{(}\mathbb{C}^{k},\mathbb{C}^{r}\bigr{)}$. Then $X_{+}$ and $X_{-}$ from above are GIT quotients of this representation associated to the characters
$+\colon\ (g_{1},g_{2})\mapsto\det(g_{1})\cdot\det(g_{2})\qquad\text{and}\qquad-\colon\ (g_{1},g_{2})\mapsto\det(g_{1})^{-1}\cdot\det(g_{2}),$
@@
1.2 Statement of theorem
With $X_{\pm}$ as above, consider the infinite-dimensional (symplectic) vector space
$\mathcal{H}_{X_{\pm}}=H^{*}_{\mathbb{T}}(X_{\pm})\bigl{(}\!\bigl{(}z^{-1}\bigr{)}\!\bigr{)}[\![Q,q_{\pm}]\!].$
@@ 1.2 Statement of theorem
Using the relative quasimap theory of [29], one can define $I$-functions $I_{X_{\pm}}(Q,q_{\pm},z)$ lying in $\mathcal{H}_{X_{\pm}}$. These functions are shown to lie on Givental’s Lagrangian cone $\mathcal{L}_{X_{\pm}}\subset\mathcal{H}_{X_{\pm}}$ for the respective spaces and may therefore be expressed as generating functions of genus-zero Gromov–Witten invariants of $X_{\pm}$. See Section 3 for details.
One can then extend $\mathcal{H}_{X_{\pm}}$ by ~$\widetilde{\mathcal{H}}_{X_{\pm}}=\mathcal{H}_{X_{\pm}}[\log z]\bigl{[}z^{-1/2}\bigr{]}$, and following Iritani in [22], define the map
$\Psi\colon\ K^{0}_{\mathbb{T}}(X)\to\widetilde{\mathcal{H}}_{X},\qquad A\mapsto z^{-\mu}z^{\rho}\bigl{(}\hat{\Gamma}_{X}\cup(2\pi{\rm i})^{\operatorname{deg}_{0}/2}\operatorname{ch}^{\mathbb{T}}(A)\bigr{)},$
where ${\rm i}=\sqrt{-1}$, $\mu$ is the grading operator, $\rho=c_{1}^{\mathbb{T}}(X_{\pm})$, and $\hat{\Gamma}_{X}$ is the Gamma class – a characteristic class associated with the Gamma function $\Gamma(1+z)$.
We prove the following.
@@
There exists a linear symplectic isomorphism $\mathbb{U}\colon\widetilde{\mathcal{H}}_{X_{-}}\to\widetilde{\mathcal{H}}_{X_{+}}$ such that the following diagram commutes:
${K^{0}_{\mathbb{T}}(X_{-})}$${K^{0}_{\mathbb{T}}(X_{+})}$~${\widetilde{\mathcal{H}}_{X_{-}}}$~${\widetilde{\mathcal{H}}_{X_{+}}.}$$\scriptstyle{\mathbb{FM}}$$\scriptstyle{\Psi_{-}}$$\scriptstyle{\Psi_{+}}$$\scriptstyle{\mathbb{U}}$ (1.2)
@@
Theorem 1.2 generalizes previous known cases of the crepant transformation conjecture in two directions (see Section 1.4 for more detailed comparisons with previous work). First, to our knowledge it is one of the first examples of a wall crossing result based on analytic continuation of generating functions for a variation of GIT quotients by a non-abelian group – in this case ${\rm GL}_{r}$. Second, it provides an example of the conjecture for fiber bundles over an arbitrary smooth base $B$. Thus, in the special case of the quotient by ${\rm GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over B, our result generalizes a particular case of the main result in [14].
We expect these techniques to be applicable beyond Grassmann flops, to more general variation of GIT quotients with respect to actions of ${\rm GL}_{n}$. In this way, the current paper may be viewed as a proof of concept. It would also be very interesting to extend these methods to quotients by other non-abelian groups. We hope to revisit this question in future work.
1.3 Strategy of proof
@@
$I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,\pm},\dots,q_{r,\pm},z)\qquad\text{at $q_{1,\pm}=\cdots=q_{r,\pm}=q_{\pm}$},$
@@
where $I_{X_{T,\pm}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}$ is a twist of the $I$-function for the associated abelian quotient by the root bundle for $G$, which we denote by $\mathfrak{g}/\mathfrak{t}$. See Section 3 for details.
@@
The technical heart of the proof relies on computing the analytic continuation of $I_{X_{+}}$ along the path $\hat{\gamma}$ and using this to determine $\mathbb{U}$. It is not clear how to compute $\widetilde{I_{X_{+}}}$ directly. Instead, we compute the analytic continuation of $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\dots,q_{r,+},z)$ along each of the variables $q_{k,+}$ independently. That is to say, we analytically continue $I_{X_{T,+}}^{e_{\mathbb{T}}(\mathfrak{g}/\mathfrak{t})}(Q,q_{1,+},\dots,q_{r,+},z)$ along a concatenated path $\gamma_{1}\star\cdots\star\gamma_{r}$, where each $\gamma_{k}$ leaves the variables $q_{i,+}$ constant for $i\neq k$, and goes from 00 to $\infty$ in the $q_{k,+}$ direction. We then homotope the path $\gamma_{1}\star\cdots\star\gamma_{r}$ to a path $\hat{\gamma}$ which lies entirely in the locus $q_{1,+}=\cdots=q_{r,+},$ thereby obtaining $\widetilde{I_{X_{+}}}$.
@@
Finally, we must verify that the linear map $\mathbb{U}$ identifying $I_{X_{-}}$ with $\widetilde{I_{X_{+}}}$ is compatible with the Fourier–Mukai transform. This involves an explicit computation of $\mathbb{FM}$ on generators of $K^{0}_{\mathbb{T}}(X_{-})$. The proof that the map $\mathbb{U}$ in (1.2) satisfies (1.3) then involves a nontrivial identity of anti-symmetric functions (Lemma 5.2).
1.4 Relation to other work
@@
The structure of the proof follows that given in [10, 11, 14, 16]. In particular, the formulation of the correspondence and the proof via localization is closely related to [14]. The new ingredients in the current article are the homotopy of the path of analytic continuation, the generalization to the relative setting, and the combinatorial identity verifying agreement with $\mathbb{FM}$, which does not appear in the case of abelian variation of GIT.
In the special case of the quotient by ${\rm GL}_{1}=\mathbb{C}^{*}$, where the two GIT quotients are toric bundles over $B$, (a compactification of) this example was studied as the local model in [26] and a similar result was obtained at the level of $D$-modules and quantum cohomology using different methods. It is possible that the comparison of $I$-functions in this paper may be obtained from their results, although it is not immediately apparent how to pursue such a strategy. On the other hand, our comparison with the $K$-theoretic Fourier–Mukai transform appears to be new even in this case.
The Gromov–Witten theory of spaces related by variation of GIT quotients by nonabelian groups has also been studied in [21] in the very general setting of GIT quotients of a projective variety by a complex reductive group. The setting of the current paper is different, as our GIT quotients are of quasi-projective varieties, and the form of the comparison is also different. The generating functions being compared in loc. cit. are different than in this paper, and the difference between the two generating functions is expressed as an explicit sum of residues rather than via analytic continuation.
@@
In the final stages of writing this paper, we learned of a related result by Lutz–Shafi–Webb (see [28]). While both papers use the idea of homotoping the path of analytic continuation, there are significant differences in the two results and their proofs. The most important such difference is in the analysis of the symplectic transformation $\mathbb{U}$. In this paper, we compare the map $\mathbb{U}$ to the Fourier–Mukai transform $\mathbb{FM}$. In [28], they instead compare $\mathbb{U}$ with the symplectic transformation of the associated abelian wall crossing. This leads to two different proofs that the map $\mathbb{U}$ is a symplectic isomorphism and has a well-defined non-equivariant limit. See Remark 5.5 for an explicit comparison of approaches. The two papers are in fact complementary, in that our respective results may be combined to show that the abelian/non-abelian correspondence is compatible with the Fourier–Mukai transform in $K$-theory. Such a compatibility is not a priori obvious, as it is not clear how best to relate the kernel $\tilde{X}$ of the Fourier–Mukai transform of the Grassmann flop to the kernel of Fourier–Mukai transform of the associated abelian wall crossing. We hope to revisit this issue in the future to obtain a more direct and conceptual understanding of this compatibility.
2 Geometric setup